# Data-Driven LQR with Finite-Time Experiments via Extremum-Seeking Policy Iteration

### v2

@@ Abstract
In this paper, we address Linear Quadratic Regulator (LQR) problems through a novel iterative algorithm named EXtremum-seeking Policy iteration LQR (EXP-LQR). The peculiarity of EXP-LQR is that it only needs access to a truncated approximation of the infinite-horizon cost associated to a given policy. Hence, EXP-LQR does not need the direct knowledge of neither the system matrices, cost matrices, and state measurements. In particular, at each iteration, EXP-LQR refines the maintained policy using a truncated LQR cost retrieved by performing finite-time virtual or real experiments in which a perturbed version of the current policy is employed. Such a perturbation is done according to an extremum-seeking mechanism and makes the overall algorithm a time-varying nonlinear system. By using a Lyapunov-based approach exploiting averaging theory, we show that EXP-LQR exponentially converges to an arbitrarily small neighborhood of the optimal gain matrix. We corroborate the theoretical results with numerical simulations involving the control of an induction motor.
@@ I Introduction
The main contribution of this paper is the development of EXtremum-seeking Policy iteration LQR (EXP-LQR), namely, a novel data-driven strategy for solving LQR problems. Our approach does not need direct knowledge of system matrices, cost matrices, and state measurements and situates itself at the intersection of off-policy and on-policy methods. More specifically, our method only needs a finite-time truncated version of the infinite-horizon cost (obtained, e.g., by running the real system or a simulator) computed by using a suitably perturbed version of the current policy maintained by the algorithm, see the schematic representation provided in Fig. 1.
@@
Using this information, EXP-LQR iteratively improves the policy taking on an extremum-seeking mechanism and a suitable reformulation of the LQR problem. This results in an overall algorithm that we interpret as a nonlinear time-varying system, which we then analyze by using system-theoretic tools based on the so-called averaging approach (see, for example, [44, Ch. 10] and [45] for the continuous-time case or [46] for the discrete-time one). Indeed, as customary in the context of averaging theory, we focus on the so-called averaged system associated to the algorithm. In particular, the averaged system reads as a policy gradient method perturbed by errors arising from the use of the truncated cost instead of the infinite-horizon one, as well as from the derivative-free gradient approximation. More in detail, we employ a Lyapunov-based approach to ensure that the averaged system trajectories exponentially converge to an arbitrarily small neighborhood of the optimal gain matrix. Then, we use this preparatory result to achieve the same property on the trajectories of the original time-varying algorithm. This final step is supported by Theorem 1, introduced in Section II, which presents averaging-related stability results for generic discrete-time systems. To the best of the authors’ knowledge, Theorem 1 also represents a per se contribution of this work. A conference version of this paper appeared in [47]. However, in that preliminary version, the algorithm relies on oracles providing the exact infinite-horizon cost associated to the tentative gain, making it impractical for real-world scenarios where only finite-time virtual or real experiments are feasible. Moreover, certain proofs were omitted. Finally, this work includes a concrete application example involving the control of an induction motor.
@@ II Preliminaries: Averaging Theory for Discrete-Time Systems
The next assumption characterizes the convergence properties of the so-called averaged system associated to (1), i.e., the auxiliary time-invariant dynamics described by
$\chi^{{k+1}}_{\textsc{av}}=\chi^{k}_{\textsc{av}}+\gamma f_{\textsc{av}}(\chi^{k}_{\textsc{av}})\qquad\chi^{0}_{\textsc{av}}=\chi_{0},$ (3)
with $\chi^{k}_{\textsc{av}}\in\mathbb{R}^{n}$.
@@
There exists $V:\mathcal{X}\to\mathbb{R}_{+}$ such that $V\in\mathcal{C}^{1}$ and, for any $c_{0}>0$ such that $\Omega_{c_{0}}\subseteq\mathcal{X}$ and any $\rho_{\textsc{av}}\in(0,c_{0})$, there exist $\bar{\gamma}_{1},a>0$ such that, for all $\chi_{0}\in\Omega_{c_{0}}$ and for all $\gamma\in(0,\bar{\gamma_{1}})$, it holds
$V(\chi^{k}_{\textsc{av}})\leq V(\chi_{0})\exp(-\gamma ak),$ (4)
along the trajectories of (3) and as long as $V(\chi^{k})\geq\rho_{\textsc{av}}$. $\blacksquare$
@@
Consider system (1) and let Assumptions 1-3 hold. Then, for any $c_{0},c_{1}>0$ such that $\Omega_{c_{0}}\subset\Omega_{c_{1}}\subseteq\mathcal{X}$ and $\rho\in(0,c_{0})$ there exist $\bar{\gamma}>0$ and $\bar{k}\in\mathbb{N}$ such that, for all $\chi_{0}\in\Omega_{c_{0}}$ and $\gamma\in(0,\bar{\gamma})$, it holds
$V(\chi^{k})\leq c_{1},$ (5)
for all $k\in\mathbb{N}$ and
$V(\chi^{k})\leq\rho,$ (6)
for all $k\geq\bar{k}$. Moreover, the convergence to $\Omega_{\rho}$ is exponentially fast. $\blacksquare$
@@ III-A Data-Driven LQR Problem Setup
subj. to $\>x_{{t+1}}=Ax_{t}+Bu_{t},\qquad\qquad x^{0}\sim\mathcal{X}^{0},$ (7b)
where $x_{t}\in\mathbb{R}^{n}$ and $u_{t}\in\mathbb{R}^{m}$ denote, respectively, the state and the input of the system at time $t\in\mathbb{N}$, $A\in\mathbb{R}^{{n}\times{n}}$ and $B\in\mathbb{R}^{{n}\times{m}}$ represent the state and the input matrices, while $Q\in\mathbb{R}^{{n}\times{n}}$ and $R\in\mathbb{R}^{{m}\times{m}}$ are the cost matrices. As for the initial condition $x^{0}\in\mathbb{R}^{n}$, we assume that it is drawn from the uniform probability distribution $\mathcal{X}^{0}$ over the unitary-sphere. The operator $\mathbb{E}[\cdot]$ denotes the expected value with respect to $\mathcal{X}^{0}$. We require the following properties on the pairs $(A,B)$ and $(Q,R)$.
@@ III-B Model-based Gradient Method for LQR
$J(K):=\tfrac{1}{2}\operatorname{Tr}\sum_{t=0}^{\infty}(A+BK)^{t,\top}(Q+K^{\top}RK)(A+BK)^{t}.$
It is worth noting that since $x^{0}\sim\mathcal{X}^{0}$ (see problem (7)) and $\mathcal{X}^{0}$ is a uniform distribution over the unitary-radius sphere, then the set of stabilizing gains $\mathcal{K}$ coincides with the domain of the cost function $J$ [24]. Moreover, being the set $\mathcal{K}$ open [49, Lemma IV.3] and connected [49, Lemma IV.6], one could use the gradient descent method to iteratively solve problem (8) (see, e.g., [24]). Namely, at each iteration $k\in\mathbb{N}$, an estimate $K^{k}\in\mathbb{R}^{{n}\times{m}}$ of the optimal gain $K^{\star}$ could be maintained and iteratively updated according to
@@
Therefore, in our setup, it is not possible to compute $G(K^{k})$ and implement (9) because its computation would require the knowledge of the pairs $(A,B)$ and $(Q,R)$ that are both not available (cf. Assumption 5). However, for a given gain $K$ (e.g., the current estimate about the optimal gain $K^{\star}$), we assume the presence of an oracle providing the associated finite-horizon cost
$J_{T}(K):=\tfrac{1}{2}\operatorname{Tr}\sum_{t=0}^{T-1}(A+BK)^{t,\top}(Q+K^{\top}RK)(A+BK)^{t},$
where the number of samples $T\in\mathbb{N}$ represents an algorithm parameter that will be designed later. Differently from the entire cost $J(K)$ whose exact computation would require virtual or real experiments over infinite time horizons, we remark that $J_{T}(K)$ may be retrieved with finite-time virtual or real experiments using the control law $u_{t}=Kx_{t}$.
@@
We emphasize that the LQR problem (7) depends on the expectation over the system initial state $x^{0}$. However, since the latter is drawn from the uniform distribution $\mathcal{X}^{0}$ over the unitary sphere, the truncated cost $J_{T}(K)$ can be computed through ${n}$ virtual or real experiments each composed of $T$ samples of system (7b) controlled via $u_{t}=Kx_{t}$. In particular, the $i$-th experiment must be executed with $x^{0}=e_{i}$, where $e_{i}\in\mathbb{R}^{{n}}$ denotes the $i$-th vector of the canonical basis in $\mathbb{R}^{{n}}$. We also remark that retrieving $J_{T}(K)$ does not require measuring $x_{t}$, In particular, it can be obtained either measuring the output $y_{t}=x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}$ for $t\in\{0,\dots,T-1\}$ or having an oracle that directly provides the cumulative truncated cost $J_{T}(K)$ at the end of the experimental phase. $\blacksquare$
@@ IV EXP-LQR: Algorithm Description and Convergence Properties
     Experiment phase
@@
     Optimization phase:
@@
where $\mathbf{k}_{ij,\text{prd}},\phi_{ij}\in\mathbb{R}$ are the period and the phase of component $(i,j)$, respectively, for all $(i,j)\in\{1,\ldots,{n}\}\times\{1,\ldots,{m}\}$. Such a perturbed policy is used to implement the feedback control law $u_{t}=(K^{k}+\delta\mathrm{D}^{k})x_{t}$ and retrieve the corresponding finite-horizon cost $J_{T}(K^{k}+\delta\mathrm{D}^{k})$ providing an approximation of the infinite-horizon one $J(K+\delta\mathrm{D}^{k})$. This scenario may occur, for example, when a simulator of a complex system is available, but the analytical knowledge of the dynamics being implemented for the simulations is unavailable. Hence, the finite-time truncation turns out to be crucial in avoiding experiments over infinite time horizons. With $J_{T}(K^{k}+\delta\mathrm{D}^{k})$ at hand, we perform the algorithm iteration detailed in (10). Specifically, the variable $\mathrm{z}^{k}\in\mathbb{R}$ filters the variation of $J_{T}(K^{k}+\delta\mathrm{D}^{k})$ (see its update (10a)), while the evolution of the gain matrix $K^{k}$ follows the extremum-seeking update (10b).
A block diagram representation that graphically describes EXP-LQR is provided in Fig. 2.
@@
Let $\mathbf{k}_{\text{prd}}\in\mathbb{N}$ be the least common multiple of all periods $\mathbf{k}_{11,\text{prd}},\dots,\mathbf{k}_{{m}{n},\text{prd}}$. Then, it holds
@@
Now, we are in the position to provide the main result of the paper, i.e., the convergence properties of EXP-LQR.
@@
Consider EXP-LQR and let Assumptions 45, and 6 hold. Then, for any $r>0$ and $(\mathrm{z}^{0},K^{0})\in\mathbb{R}\times\mathcal{K}$, there exist $\bar{\gamma},\bar{\delta}>0$ and $\bar{k},\bar{T}\in\mathbb{N}$, such that, for all $\gamma\in(0,\bar{\gamma})$, $\delta\in(0,\bar{\delta})$, $T\geq\bar{T}$, the trajectories of (10) are bounded and it holds
$\left\|K^{k}-K^{\star}\right\|\leq r,$ (12)
for all $k\geq\bar{k}$. Moreover, the convergence to the set $\{(\mathrm{z},K)\in\mathbb{R}\times\mathcal{K}\mid\left\|K-K^{\star}\right\|\leq r\}$ is exponentially fast. $\blacksquare$

### v3

@@ Abstract
In this paper, we address Linear Quadratic Regulator (LQR) problems through a novel iterative algorithm named EXtremum-seeking Policy iteration LQR (EXP-LQR). The peculiarity of EXP-LQR is that it only needs access to a truncated approximation of the infinite-horizon cost associated to a given policy. Hence, EXP-LQR does not need the direct knowledge of neither the system and cost matrices. In particular, at each iteration, EXP-LQR refines the maintained policy using a truncated LQR cost retrieved by performing finite-time virtual or real experiments in which a perturbed version of the current policy is employed. Such a perturbation is done according to an extremum-seeking mechanism and makes the overall algorithm a time-varying nonlinear system. By using a Lyapunov-based approach exploiting averaging theory, we show that EXP-LQR exponentially converges to an arbitrarily small neighborhood of the optimal gain matrix. We corroborate the theoretical results with numerical simulations involving the control of an induction motor.
@@ I Introduction
The main contribution of this paper is the development of EXtremum-seeking Policy iteration LQR (EXP-LQR), namely, a novel data-driven strategy for solving LQR problems. Our approach does not need direct knowledge of system and cost matrices. More specifically, our method only needs a finite-time truncated version of the infinite-horizon cost (obtained, e.g., by running the real system or a simulator) computed by using a suitably perturbed version of the current policy maintained by the algorithm, see the schematic representation provided in Fig. 1.
@@
Using this information, EXP-LQR iteratively improves the policy taking on an extremum-seeking mechanism and a suitable reformulation of the LQR problem. Building on the extremum-seeking framework, our mechanism employs deterministic dither matrices to perturb the current gain and leverages a low-pass filter to enhance the gradient approximation. The design of the dither matrix, along with the incorporation of the low-pass filter and an ad-hoc policy evaluation procedure, represents a key novelty and advantage of our method compared to existing schemes in the literature, which are mostly based on random perturbations and do not exploit any filtering. We interpret the overall algorithm as a nonlinear time-varying system, which we then analyze by using system-theoretic tools based on the so-called averaging approach (see, for example, [44, Ch. 10] and [45] for the continuous-time case or [46] for the discrete-time one). Indeed, as customary in the context of averaging theory, we focus on the so-called averaged system associated to the algorithm. In particular, the averaged system reads as a policy gradient method perturbed by errors arising from the use of the truncated cost instead of the infinite-horizon one, as well as from the derivative-free gradient approximation. More in detail, we employ a Lyapunov-based approach to ensure that the averaged system trajectories exponentially converge to an arbitrarily small neighborhood of the optimal gain matrix. Then, we use this preparatory result to achieve the same property on the trajectories of the original time-varying algorithm. This final step is supported by Theorem 1, introduced in Section II, which presents averaging-related stability results for generic discrete-time systems. To the best of the authors’ knowledge, Theorem 1 also represents a per se contribution of this work. A conference version of this paper appeared in [47]. However, in that preliminary version, the algorithm relies on oracles providing the exact infinite-horizon cost associated to the tentative gain, making it impractical for real-world scenarios where only finite-time virtual or real experiments are feasible. Moreover, certain proofs were omitted. Finally, this work includes a concrete application example involving the control of an induction motor.
@@ II Preliminaries: Averaging Theory for Discrete-Time Systems
The next assumption characterizes the convergence properties of the so-called averaged system associated to (1), i.e., the auxiliary time-invariant dynamics of $\chi^{k}_{\textsc{av}}\in\mathbb{R}^{n}$ described by
$\chi^{{k+1}}_{\textsc{av}}=\chi^{k}_{\textsc{av}}+\gamma f_{\textsc{av}}(\chi^{k}_{\textsc{av}})\qquad\chi^{0}_{\textsc{av}}=\chi_{0}.$ (3)
To this end, we first introduce a continuously differentiable function $V:\mathcal{X}\to\mathbb{R}_{+}$ and, given any $c>0$, its level set defined as $\Omega_{c}:=\{x\in\mathbb{R}^{n}\mid V(x)\leq c\}$.
@@
For all $c_{0}>0$ and all $\rho_{\textsc{av}}\in(0,c_{0})$, there exist $\bar{\gamma}_{1}>0$ and $a\in(0,1)$ such that, for all $\chi_{0}\in\Omega_{c_{0}}$ and $\gamma\in(0,\bar{\gamma_{1}})$, it holds
$V(\chi^{k}_{\textsc{av}})\leq(1-\gamma a)^{k}V(\chi_{0})+\rho_{\textsc{av}},$ (4)
along the trajectories of (3) for all $k\in\mathbb{N}$. $\blacksquare$
@@
Consider system (1) and let Assumptions 1-3 hold. Then, for all $c_{0},c_{1}>0$ such that $\Omega_{c_{0}}\subset\Omega_{c_{1}}\subseteq\mathcal{X}$ and $\rho\in(0,c_{0})$, there exists $\bar{\gamma}>0$ such that, for all $\chi_{0}\in\Omega_{c_{0}}$ and $\gamma\in(0,\bar{\gamma})$, it holds
$V(\chi^{k})$ $\leq c_{1},$ (5a)
$V(\chi^{k})$ $\leq(1-\gamma a)^{k}V(\chi_{0})+\rho,$ (5b)
for all $k\in\mathbb{N}$. $\blacksquare$
@@ III-A Data-Driven LQR Problem Setup
subj. to $\>x_{{t+1}}=Ax_{t}+Bu_{t},\qquad\qquad x_{0}\sim\mathcal{X}_{0},$ (6b)
where $x_{t}\in\mathbb{R}^{n}$ and $u_{t}\in\mathbb{R}^{m}$ denote, respectively, the state and the input of the system at time $t\in\mathbb{N}$, $A\in\mathbb{R}^{{n}\times{n}}$ and $B\in\mathbb{R}^{{n}\times{m}}$ represent the state and the input matrices, while $Q\in\mathbb{R}^{{n}\times{n}}$ and $R\in\mathbb{R}^{{m}\times{m}}$ are the cost matrices. As for the initial condition $x_{0}\in\mathbb{R}^{n}$, we assume that it is drawn from the uniform probability distribution $\mathcal{X}_{0}$ over the unitary-sphere. The operator $\mathbb{E}[\cdot]$ denotes the expected value with respect to $\mathcal{X}_{0}$. We require the following properties on the pairs $(A,B)$ and $(Q,R)$.
@@ III-B Model-based Gradient Method for LQR
$J(K)\!:=\!\frac{1}{2{n}}\!\operatorname{Tr}\!\bigg{(}\!\sum_{t=0}^{\infty}(A+BK)^{t,\top}(Q\!+\!K^{\top}RK)(A+BK)^{t}\!\bigg{)}\!.$
It is worth noting that since $x_{0}\sim\mathcal{X}_{0}$ (see problem (6)) and $\mathcal{X}_{0}$ is a uniform distribution over the unitary-radius sphere, then the set of stabilizing gains $\mathcal{K}$ coincides with the domain of the cost function $J$ [24]. Moreover, being the set $\mathcal{K}$ open [49, Lemma IV.3] and connected [49, Lemma IV.6], one could use the gradient descent method to iteratively solve problem (7) (see, e.g., [24]). Namely, at each iteration $k\in\mathbb{N}$, an estimate $K^{k}\in\mathbb{R}^{{n}\times{m}}$ of the optimal gain $K^{\star}$ could be maintained and iteratively updated according to
@@
Hence, in our setup, it is not possible to compute $G(K^{k})$ and implement (8) because its computation would require the knowledge of the pairs $(A,B)$ and $(Q,R)$ that are both not available (cf. Assumption 5). However, for a given gain $K$ (e.g., the current estimate about the optimal gain $K^{\star}$), we assume the presence of an oracle providing the finite-horizon cost
$J_{T}(K)\!:=\!\frac{1}{2{n}}\!\operatorname{Tr}\!\left(\sum_{t=0}^{T-1}(A\!+\!BK)^{t,\top}(Q\!+\!K^{\top}RK)(A\!+\!BK)^{t}\!\right)\!\!,$
where the number of samples $T\in\mathbb{N}$ represents an algorithm parameter that will be designed later. Differently from the entire cost $J(K)$ whose exact computation would require virtual or real experiments over infinite-time horizons, we remark that $J_{T}(K)$ may be retrieved with finite-time virtual or real experiments using the control law $u_{t}=Kx_{t}$. Since the initial condition $x_{0}$ is drawn from the uniform distribution $\mathcal{X}_{0}$ over the unitary sphere, $J_{T}(K)$ can be exactly computed as the mean truncated cost achieved over ${n}$ experiments, namely
$J_{T}(K)=\frac{1}{{n}}\sum_{i=1}^{{n}}\mathcal{V}_{T}(K,e_{i}),$ (9)
where $\mathcal{V}_{T}(K,e_{i})$ is the truncated LQR cost obtained by running system (6b) with control gain $K$ and initial condition $e_{i}$, i.e., the $i$-th canonical basis vector, namely
$\mathcal{V}_{{t+1}}(K,e_{i})$ $=\mathcal{V}_{t}(K,e_{i})+\frac{1}{2}x_{t}^{\top}\left(Q+K^{\top}RK\right)x_{t}$ (10a)
$x_{{t+1}}$ $=(A+BK)x_{t},$ (10b)
for all $t\in\{0,\dots,T-1\}$ with $\mathcal{V}_{0}(K,e_{i})=0$ and $x_{0}=e_{i}$.
@@
Notice that the availability of the experimental truncated costs $\mathcal{V}_{T}(K,e_{1}),\dots,\mathcal{V}_{T}(K,e_{{n}})$ does not necessarily imply the knowledge of $(Q,R)$. For instance, the experimental costs $\mathcal{V}_{T}(K,e_{i})$ in (10) may be obtained from sensor data or users’ feedback, as in the personalized optimization framework [50]. We also note that the unavailability of $(Q,R)$ renders methods relying on a learning phase for $(A,B)$ only (see, e.g., [34, 35, 36]) inapplicable. $\blacksquare$
@@ IV EXP-LQR: Algorithm Description and Convergence Properties
     Experiment phase (policy evaluation)
@@
     Optimization phase (policy improvement)
@@
where $\mathbf{k}_{ij,\text{prd}}\in\mathbb{Q}$ and $\phi_{ij}\in\mathbb{R}$ are the period and the phase of component $(i,j)$, respectively, for all $(i,j)\in\{1,\ldots,{n}\}\times\{1,\ldots,{m}\}$. Such a perturbed policy is used to implement the feedback control law $u_{t}=(K^{k}+\delta\mathrm{D}^{k})x_{t}$ and retrieve the corresponding finite-horizon cost $J_{T}(K^{k}+\delta\mathrm{D}^{k})$ providing an approximation of the infinite-horizon one $J(K+\delta\mathrm{D}^{k})$. This scenario may occur, for example, when a simulator of a complex system is available, but the analytical knowledge of the dynamics being implemented for the simulations is unavailable. Hence, the finite-time truncation turns out to be crucial in avoiding experiments over infinite time horizons. With $J_{T}(K^{k}+\delta\mathrm{D}^{k})$ at hand, we perform the algorithm iteration detailed in (11). Specifically, the variable $\mathrm{z}^{k}\in\mathbb{R}$ filters the variation of $J_{T}(K^{k}+\delta\mathrm{D}^{k})$ (see its update (11a)), while the evolution of the gain matrix $K^{k}$ follows the extremum-seeking update (11b). Unlike most existing algorithms (see, e.g., [23]) that rely on random perturbations, our extremum-seeking-based scheme benefits from the use of a deterministic dither matrix $\mathrm{D}^{k}$, whose design conditions will be detailed in the following. The combination of the dither matrix, the low-pass filter $\mathrm{z}^{k}$, and ad-hoc deterministic experiments (see (9) and (10)) constitute the main novelties and the key advantages of the proposed scheme with respect to the existing methods. A block diagram representation that graphically describes EXP-LQR is provided in Fig. 2.
@@
The periods $\mathbf{k}_{11,\text{prd}},\dots,\mathbf{k}_{{m}{n},\text{prd}}$ admit a least common multiple $\mathbf{k}_{\text{prd}}\in\mathbb{N}$. Moreover, it holds
@@
It is worth noting that a dither satisfying (12a) and (12b) can be interpreted as containing a set of orthogonal functions, in line with classic results from system identification theory. Moreover, we also want (12c) to be verified to improve the accuracy of the gradient estimation provided by the extremum seeking machinery. Conditions (12a)-(12c), which are in discrete time, are equivalent to those described in [40, §2.1] in continuous time. Now, we are in the position to provide the main result of the paper, i.e., the convergence properties of EXP-LQR.
@@
Consider EXP-LQR and let Assumptions 45, and 6 hold. Then, for all $r>0$ and $(\mathrm{z}^{0},K^{0})\in\mathbb{R}\times\mathcal{K}$, there exist $\bar{\gamma},\bar{\delta},a_{0}>0$, $a\in(0,1)$, and $\bar{T}\in\mathbb{N}$, such that, for all $\gamma\in(0,\bar{\gamma})$, $\delta\in(0,\bar{\delta})$, $T\geq\bar{T}$, the trajectories of (11) are bounded and satisfy
$K^{k}+\delta\mathrm{D}^{k}$ $\in\mathcal{K}$ (13a)
$\left\|K^{k}-K^{\star}\right\|$ $\leq a_{0}(1-\gamma a)^{k}+r,$ (13b)
for all $k\in\mathbb{N}$. $\blacksquare$
@@
Theorem 2 requires an initial gain $K^{0}$ that stabilizes the unknown pair $(A,B)$. We note that this does not necessarily require knowledge of $(A,B)$, see, e.g., the data-driven approach in [12], which computes stabilizing gains from system samples.
Remark 2.
Theorem 2 formally guarantees that, for sufficiently small values of $\gamma$ and $\delta$, and sufficiently large values of $T$, EXP-LQR achieves the desired convergence properties. The proofs provide all the necessary steps to determine such values. However, in practice, obtaining them analytically may be complex. As a result, one may prefer to determine $\bar{\gamma}$, $\bar{\delta}$, and $\bar{T}$ empirically, starting from reasonable initial guesses and then iteratively decreasing $\gamma$ and $\delta$, and increasing $T$, until the desired convergence properties are achieved. $\blacksquare$